# QoS-Aware Load Balancing in the Computing Continuum via Multi-Player Bandits

### v1

@@ IV-A Assumptions
  • The system is well-provisioned: a centralized oracle allocation of weights exists that satisfies the QoS threshold $\tau$ for all clients.
@@ IV-B Objective
with $a_{k,j}$ denoting the instance chosen by load balancer $k$ for request $j$.
@@ IV-C Key Properties
  • Heterogeneous rewards: Due to per-LB network latency $l_{k,m}^{n}$.
@@ IV-D Regret Definition
Regret is the standard performance measure in multi-armed bandit (MAB) problems. It quantifies the cumulative loss incurred by a learning algorithm compared to an oracle that always selects the optimal action with full knowledge of the underlying reward distributions [17]. A low (sublinear) regret indicates that the learning policy efficiently balances exploration and exploitation, ultimately achieving performance comparable to that of the oracle. In the context of this work, each load balancer acts as a learning agent that adaptively adjusts its routing weights to maximize QoS success. Therefore, the regret represents the cumulative QoS performance gap between the decentralized routing policy and an oracle weight allocation that knows the success probabilities of all service instances.
@@ V Decentralized Algorithms for QoS-Aware Load Balancing in CC
We propose a set of decentralized algorithms for QoS-aware load balancing in the CC. Each load balancer (LB) operates as an independent learner that continuously adapts its routing weights toward the set of service instances that can meet its clients’ QoS requirements. The algorithms are designed to be fully decentralized and lightweight, relying only on local latency measurements and limited placement updates from the orchestrator. This enables scalability across large CC deployments with thousands of nodes and dynamic instance placement.
Each LB operates in continuous time indexed by discrete units $t=1,2,3,\dots$. The adaptation process is organized into discrete update intervals indexed by $d=1,2,3,\dots$, where each step spans a fixed horizon of $H_{d}$ time units. At the end of each decision step, the LB updates its estimates of instance QoS performance, recalculates routing weights, and adjusts its exploration budget. LBs receive placement updates (i.e., instance additions or removals) only for nodes within their potential QoS reach (nodes that can satisfy the latency threshold $\tau$) to minimize monitoring and communication overhead.
@@ V-A QoS Pool Maintenance
Our prior work [2] introduces QoS pool: a set of service instances capable of fulfilling the specified QoS requirements. Algorithm 1 describes the main adaptive control loop executed by each LB, which maintains the service QoS pools. At the beginning of every decision step, the LB updates its local view of the available instances by applying placement updates received from the orchestrator (line 3). The feasible set $\mathcal{F}_{k}(t)$ includes all instances within the latency bound $\tau$ that are currently not in cooldown (more on cooldown in Algorithm 2). Each instance $m\in\mathcal{F}_{k}(t)$ is then evaluated using empirical QoS estimates $\hat{\mu}_{k,m}(t)$ obtained with Kernel Density Estimation (KDE) [^4] Kernel Density Estimation (KDE) is a non-parametric technique for estimating the probability density function of an unknown distribution from observed samples. Rather than producing a blocky histogram, KDE places smooth kernel functions at each data point and aggregates them to form a continuous density estimate [18]., computed over a sliding observation window of $W$ time units, representing the probability that latency $\ell_{k,m}(t)$ remains below the target $\tau$ (line 13).
Instances with an estimated QoS success probability $\hat{\mu}_{k,m}(t)$ exceeding the target threshold $\rho$ are assigned to the exploitation pool $\mathcal{E}_{k}(t)$, while the remaining feasible instances from $\mathcal{F}_{k}(t)$ form the exploration pool $\mathcal{X}_{k}(t)$. The overall QoS pool $\mathcal{Q}_{k}(t)$ is defined as their union (line 20).
Routing weights $w_{k,m}(t)$ are computed proportionally to QoS scores within each pool. The total weight is divided between the exploration and exploitation pools according to the exploration rate $\varepsilon(t)$ (line 21). Then, per-pool scores are derived: exploitation scores increase with the excess of $\hat{\mu}_{k,m}(t)$ over the QoS target $\rho$, while exploration scores increase as $\hat{\mu}_{k,m}(t)$ approaches $\rho$ (line 22). To prevent zero-valued scores and ensure stable normalization, a small smoothing factor $\eta$ is added to each score before normalization. These scores are normalized within their respective pools, producing the final routing weights $w_{k}(t)$ (line 24), with instances outside the QoS pool assigned zero weight. As the overall QoS performance $\mathrm{QoS}_{c}(t)$ stabilizes, the exploration rate $\varepsilon(t)$ decays (line 29); if degradation is detected, exploration is reset to adapt to the current system state (line 27).
@@
21: Allocate pool budgets: $W^{\mathrm{E}}(t)\leftarrow 1-\varepsilon(t)$ , $W^{\mathrm{X}}(t)\leftarrow\varepsilon(t)$
@@
23: Normalize scores within pool budgets
@@
25: Exploration schedule:
26:if $\mathrm{QoS}_{c}(t)<\mathrm{QoS}_{c}(t-W)$ then
27:  reset $\varepsilon(t{+}1)\leftarrow 1-\rho$
28:else
29:  decay $\varepsilon(t{+}1)\leftarrow\varepsilon(t)\cdot\gamma$
30:end if
31:end for
@@ V-B Request Routing
Algorithm 2 defines the per-request routing process that operates continuously in parallel with weight maintenance. Upon receiving a request, each LB selects a target instance $m$ from its current QoS pool $\mathcal{Q}_{k}(t)$ using the Smooth Weighted Round Robin (SWRR) policy based on the weight vector $w_{k}(t)$. SWRR is an implementation of the round robin mechanism, introduced by the NGINX HTTP load balancer [19] and designed to smooth the distribution of requests over time, preventing high-weight targets from receiving large bursts of requests while others wait excessively to receive new requests. The request is routed according to the weights, while the observed latency $\ell_{k,m}(t)$ and QoS success $r_{k,m}(t)$ are stored as feedback for the next decision window (line 5). Instances that repeatedly violate the latency threshold $\tau$ exceeding a predefined error count $E_{t}$ are temporarily placed in a cooldown state for a fixed duration $\Delta_{\mathrm{cd}}$ and removed from the QoS pool, ensuring that degraded instances do not continue to receive traffic (line 8). This continuous feedback loop allows each LB to adapt routing decisions dynamically to current network and service conditions without any inter-LB coordination or shared global state.
@@
4: Select instance $m\in\mathcal{Q}_{k}(t)$ using Smooth Weighted Round Robin (SWRR) with $w_{k}(t)$
@@
8:  Set $T^{\mathrm{cd}}_{m}\leftarrow t+\Delta_{\mathrm{cd}}$; remove $m$ from $\mathcal{Q}_{k}(t)$; renormalize $w_{k,\cdot}(t)$
@@ V-C Handling Instance Additions
4:   Add $m_{\text{new}}$ to $\mathcal{Q}_{k}(t)$ with initial weight $w_{k,m_{\text{new}}}(t)\leftarrow 0$
@@ V-E Regret Upper Bound
    Stationary case: When $V_{k}(T)=0$, the environment is stationary, i.e. network latencies and client loads remain constant, and the bound is determined by the $\tilde{O}(\log T)$, as proven by Auer et al. [21] for $\varepsilon$-decay policy.
@@ VI-A Deployment as a DaemonSet
QEdgeProxy is deployed like any other Kubernetes workload, as illustrated in Fig. 2. Implemented as a lightweight Golang HTTP server [^5] Source code available at https://github.com/AIoTwin/qedgeproxy, each LB connects to the Kubernetes API server to discover the nodes and service instances in the CC. To remain aware of runtime changes, each LB subscribes to node- and service-related events. To support this, we extend the standard Kubernetes control plane components with a custom module called the MQTT Informer. Implemented in Golang, [^6] Code available at https://github.com/cilicivan/k8s-mqtt-informer, this extension provides event-driven notifications for relevant cluster changes. The MQTT Informer enables lightweight MQTT-based subscriptions to events such as instance creation, removal, or node status updates for specific services. This mechanism directly supports the assumption introduced in Section IV that load balancers are informed of instance placement changes but only for nodes within their potential QoS reach. Clients access services by sending HTTP requests to their assigned QEdgeProxy, with the service name included in the request header. Once an instance is selected, the LB forwards the request directly to the corresponding pod via kube-proxy.
@@ VII-A Experimental Setup
TABLE II: Experimental configuration for evaluating QoS-aware load balancing strategies.
@@
CC Topology 30 emulated nodes with realistic latency simulation
@@ VII-A1 CC Topology Emulation
While the underlying infrastructure consisted of only 10 nodes, we emulated a larger CC topology with 30 nodes in total via Kubernetes pods. To emulate a resource-constrained CC environment, each emulated node is assigned fixed computational resources through Kubernetes resource limits [^8] https://kubernetes.io/docs/concepts/configuration/manage-resources-containers/. Specifically, every node is provisioned with 1 CPU core and 2 GB of RAM, approximating the capabilities of a typical SBC, such as a Raspberry Pi. The network topology was constructed using a latency matrix that captures the pairwise communication delays between nodes and was kept fixed throughout the experiments. Each emulated node was mapped to a city in Europe, and the latencies were drawn from round-trip time (RTT) measurements. Specifically, we relied on the WonderNetwork global ping statistics [^9] https://wondernetwork.com/pings, which provide geographically distributed latency measurements across multiple cities all over the world. Finally, network latencies between nodes were simulated using Istio VirtualService [24], which applies configurable delays to links between load balancers and service instances using Fault Injection  [^10] https://istio.io/latest/docs/tasks/traffic-management/fault-injection/.
@@ VII-A5 Load-Balancing Strategies
The second strategy, referred to as Dec-SARSA, is based on the differential SARSA algorithm for decentralized task scheduling in fog computing, proposed in [7] (see Section II). In the original formulation, each fog node operates as an autonomous agent that learns to offload tasks to neighboring nodes in order to minimize deadline violations rather than overall latency, which is in line with our QoS requirements. Each agent receives binary rewards indicating whether the latency constraint (deadline) was met and updates its action values using the SARSA algorithm, allowing it to adaptively balance local processing load and offloading decisions without requiring centralized coordination. In our adaptation, Dec-SARSA is reformulated for the CC routing context, where each load balancer functions as an independent RL agent selecting service instances on a per-request basis and learning its routing policy using the SARSA algorithm. As defined in our problem statement in Section IV, load balancers operate in a fully decentralized manner without sharing information with other nodes, thus we approximate each instance’s load through its locally observed processing latency, which serves as an implicit indicator of congestion. The agent’s state representation combines recent processing latency statistics with network proximity (RTT) estimates, and actions correspond to the available service instances. Rewards are assigned based on whether a request meets its latency deadline, directly aligning with the QoS-driven objective of our problem statement. This adaptation preserves the decentralized, self-adaptive, and model-free characteristics of the original SARSA scheduler while making it suitable for distributed load balancing in the CC environment.
The final strategy, QEdgeProxy, is our proposed solution described in Section V which we configured using an $\varepsilon$-decay factor ($\gamma=0.01$), a smoothing factor ($\eta=0.01$), an error count threshold ($E_{t}=5$), and a cooldown duration ($\Delta_{\mathrm{cd}}=10$ s). All strategies were implemented in Golang [^11] Source code available at: https://github.com/AIoTwin/qedgeproxy and deployed as LBs in the same manner as our Kubernetes-native implementation (see Section VI). This ensures that the comparison of different strategies is fair, as the only factor influencing performance is the routing strategy itself.
@@ VII-B Aggregated Results Across Topologies
VII-B1 Client-Level QoS Success Rate
Fig. 3 reports the percentage of clients whose QoS requirements are met, aggregated across the five topologies. Across all scenarios, proxy-mity exhibits poor QoS satisfaction, with proxy-mity 0.9 performing worst and proxy-mity 1.0 only marginally better, typically satisfying QoS for fewer than 50% of clients. This behavior results from routing decisions based solely on network proximity, which overloads popular instances and ignores explicit QoS constraints. Dec-SARSA substantially improves performance, satisfying QoS for approximately 70–90% of clients by incorporating latency deadlines into its reward function. However, it remains relatively slow in terms of adapting to overloads, as the Q-values evolve incrementally through differential updates and thus reflect historical performance more than short-term conditions, particularly for load balancers without local instances. In contrast, QEdgeProxy consistently achieves the highest QoS satisfaction (95–100% of clients) across all scenarios. By assigning dynamic routing weights based on estimated QoS success probabilities and balancing traffic across a pool of feasible instances using Smooth Weighted Round Robin, it adapts more rapidly and stably to changing network and load conditions.
Figure 3: Client QoS success rate aggregated across five scenarios.

### v2

@@ IV-A Assumptions
  • The system is well-provisioned: there exists a centralized oracle allocation of weights that satisfies the QoS threshold $\tau$ for all clients.
@@ IV-B Objective
with $a_{k,j}$ denoting the instance chosen by LB $k$ for request $j$.
@@ IV-C Key Properties
  • Heterogeneous rewards: Rewards are heterogeneous due to per-LB network latency $l_{k,m}^{n}$.
@@ IV-D Regret Definition
Regret is the standard performance measure in MAB problems. It quantifies the cumulative loss incurred by a learning algorithm compared to an oracle that always selects the optimal action with full knowledge of the underlying reward distributions [3]. A low (sublinear) regret indicates that the learning policy efficiently balances exploration and exploitation, ultimately achieving performance comparable to that of the oracle. In the context of this work, each LB acts as a learning agent that adaptively adjusts its routing weights to maximize QoS success. Therefore, regret represents the cumulative QoS performance gap between the decentralized routing policy and an oracle weight allocation that knows the success probabilities of all service instances.
@@ V Decentralized Algorithms for QoS-Aware Load Balancing in CC
We propose a set of decentralized algorithms for QoS-aware load balancing in the CC. Each LB operates as an independent learner, continuously adapting its routing weights toward the set of service instances that can meet its clients’ QoS requirements. The algorithms are designed to be fully decentralized and lightweight, relying only on local latency measurements and limited placement updates from the orchestrator. This enables scalability across large CC deployments with thousands of nodes and dynamic instance placement.
Each LB operates in continuous time indexed by discrete units $t=1,2,3,\dots$. The adaptation process is organized into discrete update intervals indexed by $d=1,2,3,\dots$, with each step spanning a fixed horizon of $H_{d}$ time units. At the end of each decision step, an LB updates its estimates of instance QoS performance, recalculates routing weights, and adjusts its exploration budget. LBs receive placement updates (i.e., instance additions or removals) only for nodes within their potential QoS reach (nodes that can satisfy the latency threshold $\tau$) to minimize monitoring and communication overhead.
@@ V-A QoS Pool Maintenance
Algorithm 1 describes the main adaptive control loop executed by each LB, which maintains the service QoS pools. At the beginning of every decision step, the LB updates its local view of the available instances by applying placement updates received from the orchestrator (line 3). The feasible set $\mathcal{F}_{k}(t)$ includes all instances within the latency bound $\tau$ that are currently not in cooldown (more on cooldown in Algorithm 2). Each instance $m\in\mathcal{F}_{k}(t)$ is then evaluated using empirical QoS estimates $\hat{\mu}_{k,m}(t)$ obtained with Kernel Density Estimation (KDE), [^4] Kernel Density Estimation (KDE) is a non-parametric technique for estimating the probability density function of an unknown distribution from observed samples. Rather than producing a blocky histogram, KDE places smooth kernel functions at each data point and aggregates them to form a continuous density estimate [18]. computed over a sliding observation window of $W$ time units, representing the probability that latency $\ell_{k,m}(t)$ remains below the target $\tau$ (line 13).
Instances with an estimated QoS success probability $\hat{\mu}_{k,m}(t)$ exceeding the target success ratio $\rho$ are assigned to the exploitation pool $\mathcal{E}_{k}(t)$, while the remaining feasible instances from $\mathcal{F}_{k}(t)$ form the exploration pool $\mathcal{X}_{k}(t)$. The overall QoS pool $\mathcal{Q}_{k}(t)$ is defined as their union (line 20).
Routing weights $w_{k,m}(t)$ are computed proportionally to QoS scores within each pool. The total weight is divided between the exploration and exploitation pools according to the exploration rate $\varepsilon(t)$ (line 21). Then, per-pool scores are derived: exploitation scores increase with the excess of $\hat{\mu}_{k,m}(t)$ over the QoS target $\rho$, while exploration scores increase as $\hat{\mu}_{k,m}(t)$ approaches $\rho$ (line 22). To prevent zero-valued scores and ensure stable normalization, a small smoothing factor $\eta$ is added to each score before normalization. These scores are normalized within their respective pools, producing the final routing weights $w_{k}(t)$ (line 24), with instances outside the QoS pool assigned zero weight. As the measured QoS success ratio $\hat{\rho}_{k}(t)$ stabilizes, the exploration rate $\varepsilon(t)$ decays (line 28); if degradation is detected, exploration is reset to adapt to the current system state (line 26).
@@
21: Allocate pool budgets:
$W^{\mathrm{E}}(t)\leftarrow 1-\varepsilon(t),W^{\mathrm{X}}(t)\leftarrow\varepsilon(t)$
@@
23: Normalize scores within pool budgets:
$w^{\mathrm{E}}_{k,m}(t)=W^{\mathrm{E}}(t)\,\frac{s^{\mathrm{E}}_{k,m}(t)}{\sum_{j\in\mathcal{E}_{k}(t)}s^{\mathrm{E}}_{k,j}(t)},$
$w^{\mathrm{X}}_{k,m}(t)=W^{\mathrm{X}}(t)\,\frac{s^{\mathrm{X}}_{k,m}(t)}{\sum_{j\in\mathcal{X}_{k}(t)}s^{\mathrm{X}}_{k,j}(t)}$
@@
25:if $\hat{\rho}_{k}(t)<\hat{\rho}_{k}(t-W)$ then
26:  reset $\varepsilon(t{+}1)\leftarrow 1-\rho$
27:else
28:  decay $\varepsilon(t{+}1)\leftarrow\varepsilon(t)\cdot\gamma$
29:end if
30:end for
@@ V-B Request Routing
Algorithm 2 defines the per-request routing process that operates continuously in parallel with weight maintenance. Upon receiving a request, each LB selects a target instance $m$ from its current QoS pool $\mathcal{Q}_{k}(t)$ using the Smooth Weighted Round Robin (SWRR) [^5] SWRR is an implementation of the round robin mechanism, introduced by the NGINX HTTP load balancer [19] and designed to smooth the distribution of requests over time, preventing high-weight targets from receiving large bursts of requests while others wait excessively to receive new requests. policy based on the weight vector $w_{k}(t)$. The request is routed according to the weights, while the observed latency $\ell_{k,m}(t)$ and QoS success $r_{k,m}(t)$ are stored as feedback for the next decision window (line 5). Instances that repeatedly violate the latency threshold $\tau$ exceeding a predefined error count $E_{t}$ are temporarily placed in a cooldown state for a fixed duration $\Delta_{\mathrm{cd}}$ and removed from the QoS pool, ensuring that degraded instances do not continue to receive traffic (line 8). This continuous feedback loop allows each LB to adapt routing decisions dynamically to current network and service conditions without any inter-LB coordination or shared global state.
@@
4: Select instance $m\in\mathcal{Q}_{k}(t)$ using SWRR with $w_{k}(t)$
@@
8:  Set $T^{\mathrm{cd}}_{m}\leftarrow t+\Delta_{\mathrm{cd}}$; remove $m$ from $\mathcal{Q}_{k}(t)$; renormalize $w_{k}(t)$
@@ V-C Handling Instance Additions
4:   Add $m_{\text{new}}$ to $\mathcal{Q}_{k}(t)$; init. weight $w_{k,m_{\text{new}}}(t)\leftarrow 0$
@@ V-E Regret Upper Bound
    Stationary case: When $V_{k}(T)=0$, the environment is stationary, i.e. network latencies and client loads remain constant, and the bound is determined by $\tilde{O}(\log T)$, as proven by Auer et al. [21] for the $\varepsilon$-decay policy.
@@ VI-A Deployment as a DaemonSet
QEdgeProxy is deployed like any other Kubernetes workload, as illustrated in Fig. 2. Implemented as a lightweight Golang HTTP server, [^6] Source code: https://github.com/AIoTwin/qedgeproxy/tree/tsc-paper each LB connects to the Kubernetes API server to discover the nodes and service instances in the CC. To remain aware of runtime changes, each LB subscribes to node- and service-related events. To support this, we extend the standard Kubernetes control plane components with a custom module called the MQTT Informer. Implemented in Golang, [^7] Source code: https://github.com/AIoTwin/k8s-mqtt-informer this extension provides event-driven notifications for relevant cluster changes. The MQTT Informer enables lightweight MQTT-based subscriptions to events such as instance creation, removal, or node status updates for specific services. This mechanism directly supports the assumption introduced in Section IV that load balancers are informed of instance placement changes but only for nodes within their potential QoS reach. Clients access services by sending HTTP requests to their assigned QEdgeProxy, with the service name included in the request header. Once an instance is selected, the LB forwards the request directly to the corresponding pod via kube-proxy.
@@ VII-A Experimental Setup
TABLE II: Experimental configuration for evaluating QoS-aware load-balancing strategies.
@@
CC Topology 30 emulated nodes with realistic network simulation
@@ VII-A1 CC Topology Emulation
While the underlying infrastructure consisted of 10 nodes, we emulated a larger CC topology with 30 nodes in total via Kubernetes pods. To emulate a resource-constrained CC environment, each emulated node is assigned fixed computational resources through Kubernetes resource limits. [^9] https://kubernetes.io/docs/concepts/configuration/manage-resources-containers/ Specifically, every node is provisioned with 1 CPU core and 2 GB of RAM, approximating the capabilities of a typical single-board computer, such as a Raspberry Pi. The network topology was constructed using a latency matrix that captures the pairwise communication delays between nodes and was kept fixed throughout the experiments. Each emulated node was mapped to a city in Europe, and the latencies were drawn from round-trip time (RTT) measurements. Specifically, we relied on the WonderNetwork global ping statistics, [^10] https://wondernetwork.com/pings which provide geographically distributed latency measurements across multiple cities all over the world. Finally, network latencies between nodes were simulated using Istio VirtualService [24], which applies configurable delays to links between load balancers and service instances using Fault Injection. [^11] https://istio.io/latest/docs/tasks/traffic-management/fault-injection/
@@ VII-A5 Load-Balancing Strategies
The second strategy, referred to as Dec-SARSA, is based on the differential SARSA algorithm for decentralized task scheduling in fog computing, proposed in [7] (see Section II). In the original formulation, each fog node operates as an autonomous agent that learns to offload tasks to neighboring nodes in order to minimize deadline violations rather than overall latency, which is in line with our QoS requirements. Each agent receives binary rewards indicating whether the latency constraint (deadline) was met and updates its action values using the SARSA algorithm, allowing it to adaptively balance local processing load and offloading decisions without requiring centralized coordination. In our adaptation, Dec-SARSA is reformulated for the CC routing context, where each load balancer functions as an independent RL agent selecting service instances on a per-request basis and learning its routing policy using the SARSA algorithm. As defined in our problem statement in Section IV, load balancers operate in a fully decentralized manner without sharing information with other nodes, thus we approximate each instance’s load through its locally observed processing latency, which serves as an implicit indicator of congestion. The agent’s state representation combines recent processing latency statistics with network proximity (RTT) estimates, and actions correspond to the available service instances. Rewards are assigned based on whether a request meets its latency deadline, directly aligning with the QoS-driven objective of our problem statement. This adaptation preserves the decentralized, self-adaptive, and model-free characteristics of the original SARSA scheduler while making it suitable for decentralized load balancing in the CC environment.
The final strategy, QEdgeProxy, is our proposed solution described in Section V which we configured using an $\varepsilon$-decay factor ($\gamma=0.01$), a smoothing factor ($\eta=0.01$), an error count threshold ($E_{t}=5$), and a cooldown duration ($\Delta_{\mathrm{cd}}=10$ s). All strategies were implemented in Golang [^12] Source code and experimental results: https://github.com/AIoTwin/qedgeproxy/tree/tsc-paper and deployed as LBs in the same manner as our Kubernetes-native implementation (see Section VI). This ensures that the comparison of different strategies is fair, as the only factor influencing performance is the routing strategy itself.
@@ VII-B Aggregated Results Across Topologies
VII-B1 Client-Level QoS Satisfaction Rate
Fig. 3 reports the percentage of clients whose QoS requirements are met, aggregated across the five topologies. Across all scenarios, proxy-mity exhibits poor QoS satisfaction, with proxy-mity 0.9 performing worst and proxy-mity 1.0 only marginally better, typically satisfying QoS for fewer than 50% of clients. This behavior results from routing decisions based solely on network proximity, which overloads popular instances and ignores explicit QoS constraints. Dec-SARSA substantially improves performance, satisfying QoS for approximately 70–90% of clients by incorporating latency deadlines into its reward function. However, it remains relatively slow in terms of adapting to overloads, as the Q-values evolve incrementally through differential updates and thus reflect historical performance more than short-term conditions, particularly for load balancers without local instances. In contrast, QEdgeProxy consistently achieves the highest QoS satisfaction (95–100% of clients) across all scenarios. By assigning dynamic routing weights based on estimated QoS success probabilities and balancing traffic across a pool of feasible instances using SWRR, it adapts more rapidly and stably to changing network and load conditions.
Figure 3: Client QoS satisfaction rate aggregated across five scenarios.